OpenStudy (anonymous):

how much energy does an object with a mass of 5 grams have? how to plug in e=mc squared?

OpenStudy (anonymous):

M is mass which is 5.

OpenStudy (anonymous):

but what is E

OpenStudy (anonymous):

\(E = 5c^2\)

OpenStudy (dan815):

depends on the speed its travelling at

OpenStudy (dan815):

is your mass travelling at light speed thats the only time you can use e=mc^2

OpenStudy (anonymous):

lets say it is

OpenStudy (dan815):

then ya E= mc^2 u know m=5*10^-3 kg and c=3.0*10^8 m/s

OpenStudy (wolf1728):

I believe for the E=mc² equation to work, it deals with mass being converted directly to energy. Convert a paper clip (about 1 gram) directly to energy and you get the energy equivalent to the Hiroshima bomb. (about 20 kilotons)

OpenStudy (anonymous):

alright thanks guys

OpenStudy (wolf1728):

Want to learn a bit more? Here's a good link http://www.1728.org/einstein.htm (Well it is a web page of mine.)

OpenStudy (dan815):

i hate to be a stickler but kilotons is not a unit of energy

OpenStudy (dan815):

usually when they give mass they say something like 20 kilotons of TNT is that what u mean

OpenStudy (wolf1728):

It is short for the explosive power of 20 kilotons of tnt. (trinitrotoluene)

OpenStudy (dan815):

again 20 kilotons is not short for anything xD

OpenStudy (wolf1728):

Wikipedia says The "ton of TNT" is a unit of energy equal to 4.184 gigajoules Here's the link: http://en.wikipedia.org/wiki/Kiloton

OpenStudy (dan815):

but i see thanks.. good to know :)

OpenStudy (anonymous):

1.21 gigawattz! great scott

OpenStudy (wolf1728):

thanks dan815 - this message board keeps everybody sharp :-)

OpenStudy (dan815):

1.21 gigawattz is power, u need your answer in joules

OpenStudy (anonymous):

i know i was referencing a movie

OpenStudy (wolf1728):

see? good one dan815

OpenStudy (dan815):

lol

OpenStudy (wolf1728):

Back to the Future?

OpenStudy (dan815):

what movie ?

OpenStudy (dan815):

whats so great about back to the future i cant seem to sit through that movie

OpenStudy (anonymous):

the future one

OpenStudy (dan815):

that movie feels too old to watch

OpenStudy (dan815):

i mean theres some really old movies i like a lot but

OpenStudy (anonymous):

preference i presume

OpenStudy (anonymous):

depends what your into

OpenStudy (wolf1728):

1985 I believe

OpenStudy (dan815):

oh man its late for me!! i got exam tomorrow

OpenStudy (dan815):

cya guys later

OpenStudy (anonymous):

good luck

OpenStudy (wolf1728):

c ya Think I'll find another question

OpenStudy (anonymous):

bye guys

OpenStudy (wolf1728):

c ya